OpenStudy (twizttiez):

Need help! A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct. A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (twizttiez):

@waterineyes

OpenStudy (twizttiez):

Can you help?

OpenStudy (anonymous):

Line with a slope and a point given has equation of : \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

\(m\) is slope and \((x_1,y_1)\) is point given to you..

OpenStudy (twizttiez):

So that mean C and D are already two of my answers correct?

OpenStudy (twizttiez):

Means*

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

But only one from them..

OpenStudy (anonymous):

\(x_1\) = 5 and \(y_1 = 9\).. Also \(m=7\) Just plug the values in the equation I gave you above..

OpenStudy (anonymous):

\(y_1 = -9\).. Sorry..

OpenStudy (twizttiez):

y-(-9)=7(x-x^1) Right?

OpenStudy (twizttiez):

@waterineyes

OpenStudy (twizttiez):

So using that i got C,B and A

OpenStudy (anonymous):

\(x_1 = 5\)..

OpenStudy (anonymous):

C, B and A as your answers??

OpenStudy (twizttiez):

ok sorry I'm not the best at math :/

OpenStudy (twizttiez):

Thats what i got is that not right?

OpenStudy (anonymous):

\(y-(-9) = 7(x-5) \implies (y+9) = 7(x-5)\)..

OpenStudy (anonymous):

yes, C, B and A are not right, so eliminate that.. One option left..

OpenStudy (twizttiez):

So D ok

OpenStudy (anonymous):

There it is..!! :)

OpenStudy (twizttiez):

So i know i have D what else can i do?

OpenStudy (twizttiez):

Oh ok so only D is correct?

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

Misread the question, there can be more answers..

OpenStudy (anonymous):

So: you got this: \(y+9 = 7(x-5)\) now distribute 7 to the brackets..

OpenStudy (anonymous):

\[a(b+c) = ab +ac ----> Distribution \ Property\]

OpenStudy (twizttiez):

So B?

OpenStudy (anonymous):

Great..

OpenStudy (twizttiez):

Thank's so much could you help on a few more?

OpenStudy (anonymous):

you are more intelligent than me.!! :)

OpenStudy (twizttiez):

Nooo not at all xD

OpenStudy (anonymous):

One more to go still..

OpenStudy (twizttiez):

Can i open up a new question and tag you?

OpenStudy (twizttiez):

Oh one more huh? Ok 1 sec lemme look

OpenStudy (anonymous):

\(7x - y =44\) Multiply by \(-1\) both the sides..

OpenStudy (twizttiez):

That'd make it A

OpenStudy (anonymous):

Good.. :)

OpenStudy (twizttiez):

Yay could you help on a couple more?

OpenStudy (anonymous):

So, finally, except C, all are correct..!! :P

OpenStudy (anonymous):

sure, if you want, you can ask me here too..

OpenStudy (twizttiez):

You don't want medals?

OpenStudy (twizttiez):

It's only fair you're helping me and all

OpenStudy (anonymous):

No.. One is more than sufficient to me..!! :P

OpenStudy (twizttiez):

Ok then :D

OpenStudy (anonymous):

Do you know what are x and y-intercepts??

OpenStudy (twizttiez):

Yeah wound i be using rise and run?

OpenStudy (twizttiez):

Would*

OpenStudy (anonymous):

I think, in this question, if you know what is your y-intercept, you are done..

OpenStudy (twizttiez):

Then i think it would be A

OpenStudy (twizttiez):

Wait nvm it's B

OpenStudy (anonymous):

are their multiple answers to select from?

OpenStudy (anonymous):

*there.

OpenStudy (twizttiez):

And no there arean't it's only one this time

OpenStudy (anonymous):

Wait..

OpenStudy (twizttiez):

And i think it's B because the y intercept is where a line crosses the Y axis right?

OpenStudy (anonymous):

Calm down and take deep breath..:P

OpenStudy (twizttiez):

Lol sorry im kinda in a hurry xD

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

What you got for y-intercept?

OpenStudy (twizttiez):

? elaborate

OpenStudy (twizttiez):

Then again it could be C

OpenStudy (anonymous):

How you came to know line is crossing through 2 but not -2??

OpenStudy (anonymous):

I tell you, okay?

OpenStudy (twizttiez):

Good point...

OpenStudy (twizttiez):

Ok

OpenStudy (anonymous):

y-intercept: It is the point where line cuts the y-axis.. It is found by putting \(x=0\) in your equation..

OpenStudy (anonymous):

when you put \(x=0\), what will you get?

OpenStudy (twizttiez):

A 0 y intercept right?

OpenStudy (anonymous):

I mean what value of y, you will get?

OpenStudy (anonymous):

\(y = \frac{1}{3}x -2\), put \(x=0\) and find \(y\)..

OpenStudy (twizttiez):

So it would be y=1/3x-0

OpenStudy (anonymous):

Dear, put \(x=0\)..

OpenStudy (twizttiez):

Oh :/ i apologize

OpenStudy (twizttiez):

I feel a tad bit slow :/

OpenStudy (anonymous):

\[y = \frac{1}{3} \color{red}{x} -2 \implies y = \frac{1}{3} \color{red}{(0)}-2 \implies y = -2\]

OpenStudy (anonymous):

we are all learning here, so no need to apologize.. :)

OpenStudy (twizttiez):

Ohhhhh ok so looking back at my answer choices A seems right

OpenStudy (anonymous):

Can I make it a point now : as y-intercept is a point.. \((0,-2)\) implying that when x is 0, is -2.. Getting?

OpenStudy (anonymous):

*y

OpenStudy (twizttiez):

Ok that makes sense

OpenStudy (anonymous):

So, y-intercept is : \((0,-2)\).. x-intercept : It tells you where the line cuts x-axis. It is found by putting \(y=0\) in your equation..

OpenStudy (anonymous):

Can you find value of \(x\), when \(y=0\) ??

OpenStudy (twizttiez):

Hmm one moment

OpenStudy (twizttiez):

It would be -2 if im wrong im sorry :/

OpenStudy (anonymous):

\[\color{red}{y} = \frac{1}{3} x - 2 \implies \color{red}{0} = \frac{1}{3}x-2\]

OpenStudy (anonymous):

Add 2 both the sides,..

OpenStudy (anonymous):

Then multiply by 3 both sides to get \(x\)..

OpenStudy (twizttiez):

Ok so with that equation you just made i would add 2 to both sides then once i get my answer for that i would then multiply correct?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

Firstly add 2 and show me what you got..

OpenStudy (twizttiez):

2+0=2 2-2=0 1/3x+2

OpenStudy (anonymous):

If you have some doubt or difficulty, don't hesitate to tell me..

OpenStudy (twizttiez):

I have lots of it :/

OpenStudy (anonymous):

\[0 = \frac{1}{3}x -2 \\ 0 \color{green}{+2} = \frac{1}{3}x - 2 \color{green}{+2} \\ 2 = \frac{1}{3}x \cancel{-2}\cancel{+2} \\ 2 = \frac{1}{3}x\]

OpenStudy (anonymous):

Is it making sense?

OpenStudy (anonymous):

Now multiply by \(3\) both sides..